OpenStudy (anonymous):

will fan and metal just the answer and explation How many solutions does the equation 4s + 8 = 3 + 5 + 4s have? Two None One Infinitely many

OpenStudy (anonymous):

can you plz help me?

OpenStudy (welshfella):

Hint subtract 4s from both sides of the = sign and see what you get

OpenStudy (welshfella):

4s + 8 = 3 + 5 + 4s 4x + 8 = 4s + 8

OpenStudy (welshfella):

* thats should be 4s not 4x

OpenStudy (welshfella):

4x + 8 = 4s + 8 now we can put any value we like for s and the equation will still balance 4*1 + 8 = 4*1 + 8 4*2 + 8 = 4*2 + 8 and so on....

OpenStudy (welshfella):

we can put 5000006 for s and it will still balance so what do you think the answer is?

OpenStudy (anonymous):

So s is on both side correct?

OpenStudy (welshfella):

yes - I wrote x again!! its s of course

OpenStudy (welshfella):

well what do you think it is 1,2 none or infiniteley many

OpenStudy (anonymous):

one or 2

OpenStudy (welshfella):

how can it be one or 2? when i'v given you 3 already i can give you more - it can be any real number -4, -9, 60009, 1 etc etc

OpenStudy (anonymous):

Ohh okay I get it now It would be D

OpenStudy (welshfella):

right

OpenStudy (anonymous):

:D thank you

OpenStudy (welshfella):

you can any number for s and the left side of the equation will equal the right side

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

oki I fanned and metaled you :D

OpenStudy (anonymous):

bye